If a female basketball player who weighs 135 lbs, with a body fat % of 28% and is currently in her off-season doing 3-4 strength workouts per week and 2-3 conditioning sessions per week wants to reduce her body fat % to 24%, please answer the following questions:

1. What would her new body weight be if she accomplishes this goal? (assuming her fat-free mass stays the same).

2. What would her total daily energy expenditure be if the Cunningham equation was used to predict resting metabolic rate with an activity factor of 1.75?

3. What is a recommended timeline for that amount of weight loss?

4. What should her calorie intake be if she wants to lose weight during this timeline?

Cunningham Equation: Resting Metabolic Rate (kcal/d) = 370 + (21.6 x Fat-free mass [in kg])

Make sure to show all work!
Answer #1

1. Body fat loss and weight loss are two completely different concepts. Consider an instance of two individuals belonging to the same age group – an athletically built person and a normal person. Due to the extra weight of muscles, the athletic person may weigh more than his counterpart. However, the second person’s weight may be related to a higher content of body fat. Obviously, this person will be more vulnerable to weight-related issues than the athletic person. Therefore, it is important to lose body fat rather than trying to lose body weight. With new methods that measure body fat percentage, you can get an accurate report of your body fat and the amount you must lose or gain to remain fit and healthy.
So the basketball player will end up building her muscles which will thereby contribute for the overall body weight.

2.The Cunningham Equation

Step #1 Determine Resting Metabolic Rate
the trainee is 135 lbs at 28% body fat, so 0.28 equals 28% body fat

Take your weight in pounds and divide by 2.2 to get your weight in
Kilograms=135/2.2=61.36 kg

Next, take your fat percentage and multiply by your mass in kilograms,
this will give you your fat mass (FM) E.g. 61.36x0.28=17.18 kg

Subtract your (FM) from your total weight in kilograms to get your fat
free mass (FFM) E.g. 61.36-17.18=44.18 kg
Now enter your (FFM) into this formula
RMR = 500 + [22 x (FFM)] E.g. RMR=500 + [22 x 44.18]
RMR=500 + 971.96
RMR=1971.96 cal

3. Since a pound equals 3,500 calories, it's generally recommended you aim for a 500- to 1,000-calorie daily deficit to lose a pound or two per week. ... This means that in a week you could lose 5.4 pounds like in this case concerted. Much of this accelerated weight loss will come from water weight, not fat.